OpenStudy (anonymous):

multiply and simplify by factoring, if possible. assume all variables under the radicals represent nonnegitative numbers. ...........................

OpenStudy (anonymous):

\[\sqrt{2t}\sqrt{2t}\]

OpenStudy (whpalmer4):

A definition of the square root is \[\sqrt{x} * \sqrt{x} = x\] What does that suggest here?

OpenStudy (anonymous):

The answer that I keep getting is 0 but I don't know if I'm even doing it right.

OpenStudy (mathstudent55):